OpenStudy (anonymous):

What is the exact value of the expression √27+√32-√75 ?

OpenStudy (bahrom7893):

3sqrt(3) + 4sqrt(2) - 5sqrt(3) = -2sqrt(3) + 4 sqrt(2)

OpenStudy (anonymous):

3√3+4√2-5√3

OpenStudy (anonymous):

\[4\sqrt{2}-2\sqrt{3}\] in simplest radical form

OpenStudy (anonymous):

thanks you guys :)

OpenStudy (anonymous):

√27+√32-√75 = 3√3 + 4√2 - 5√3 = 4√2 - 2√3